OpenStudy (anonymous):

Preston drew a blueprint of an additional room he is planning to add to his house. On the blueprint, 1 centimeter is equal to 4 feet in the actual room. If Preston wants the room to measure 12 feet by 16 feet, what are these dimensions on the blueprint?

OpenStudy (anonymous):

A 3 cm by 4 cm B 6 cm by 8 cm C 4 cm by 6 cm

OpenStudy (anonymous):

Simple just divide the values by 4.

OpenStudy (anonymous):

divide what numbers

OpenStudy (anonymous):

4ft = 1cm then 12ft = 1/4 * 12 = 3cm and 16ft = 1/4 * 16 = 4cm